# Learning Long-Context Diffusion Policies via Past-Token Prediction

### v1

@@ 2 Related Work
Handling long sequences of high-dimensional observations has been a persistent challenge for robot learning. Many prior works mitigate it by discarding parts of the past, either through regularization strategies such as adversarial objectives [39] or information bottlenecks [28], or by summarizing trajectories using keyframes [40] or motion tracks [25]. Similarly, methods like sketch synthesis [34] and visual trace prompting [46] have been explored for generalist robot policies. These approaches often rely on assumptions about the irrelevance of specific parts of the history, which may not hold in complex tasks. In contrast, we propose a method that directly regularizes diffusion policies to retain information about past actions that would otherwise be lost from historical context.
@@ 4.1 Past-Token Prediction
We address this issue through Past-Token Prediction (PTP), an auxiliary objective that tasks the policy to predict past action tokens alongside future ones. Formally, given a sequence of observations $\mathbf{o}_{t-k:t}$, the policy is trained to jointly predict the action tokens from the past time step $t-k$ to the upcoming time step $t+l$:
$\hat{\mathbf{a}}_{t-k:t+l}=\pi_{\theta}(\mathbf{o}_{t-k:t}).$ (1)
@@ 4.3 Test-Time Verification with PTP
where each sampled candidate $\hat{\mathbf{a}}^{(i)}=(a_{t-k},\dots,a_{t+l})^{(i)}$ includes both reconstructed past actions and predicted future actions. Since the first $k-1$ actions have already been executed, we use them as a ground-truth reference and select the candidate whose reconstructed past actions best match the executed ones:
@@ Acknowledgments
We thank the members of the Stanford IRIS lab for their support and insightful discussions. We thank Joey Hejna, Moo Jin Kim, Alec Lessing, and Amber Xie for their support with the real world experiments. This work was supported in part by The Robotics and AI Institute, DARPA, and the ONR grant N00014-22-1-2621. C. Finn is a CIFAR fellow. Y. Liu is an SNSF postdoc fellow.

### v2

@@ 2 Related Work
Handling long sequences of high-dimensional observations has been a persistent challenge in robot learning. A common strategy is to reduce the input history—by discarding parts of the past via adversarial regularization [41], information bottlenecks [30], or selecting salient subsets through techniques like keyframes [42] and motion tracks [27]. Other methods construct higher-level summaries, such as sketch synthesis [36] or visual trace prompting [48], especially for generalist policies. These approaches rely on the assumption that much of the historical context is irrelevant—a simplification that may break down in temporally complex tasks. An alternative line of work attempts to model the full context in an autoregressive manner using action tokens [25, 11]. Yet, designing action tokenizers that can effectively capture long-range temporal structure remains an open problem [39]. Our method takes an orthogonal approach: we explicitly regularize diffusion policies to retain information about past actions that would otherwise be lost from historical context.
@@ 4.1 Past-Token Prediction
We address this issue through Past-Token Prediction (PTP), an auxiliary objective that tasks the policy to predict past action tokens alongside future ones. Formally, given a sequence of observations $\mathbf{o}_{t-k:t}$, the policy is trained to jointly predict the action tokens from the past time step $t-k$ to the upcoming time step $t+h$:
$\hat{\mathbf{a}}_{t-k:t+h}=\pi_{\theta}(\mathbf{o}_{t-k:t}).$ (1)
@@ 4.3 Test-Time Verification with PTP
where each sampled candidate $\hat{\mathbf{a}}^{(i)}=(a_{t-k},\dots,a_{t+h})^{(i)}$ includes both reconstructed past actions and predicted future actions. Since the first $k-1$ actions have already been executed, we use them as a ground-truth reference and select the candidate whose reconstructed past actions best match the executed ones:
@@ Acknowledgments
We thank Sergey Levine and the members of the Stanford IRIS lab for their insightful discussions. We thank Joey Hejna, Moo Jin Kim, Alec Lessing, and Amber Xie for their support with the real world experiments. This work was supported in part by The Robotics and AI Institute, DARPA, and the ONR grant N00014-22-1-2621. C. Finn is a CIFAR fellow. Y. Liu is an SNSF postdoc fellow.